OpenStudy (anonymous):

find s prime (10) for s(t)=300,000e^-.20

OpenStudy (turingtest):

s'(t)=0 if there is no t in the equation. did you write the problem correctly?

OpenStudy (anonymous):

here f prime of 10 for F(x)=300,000e^-.20

OpenStudy (turingtest):

\[f(x)=300,000e^(-0.20)\] has no variables (x's) in it, so it may as well be f(x)=5 then f'(x)=0 so f'(10)=0

OpenStudy (anonymous):

\[find s \prime (10) for s(t)=300000e^(-.20)\]

OpenStudy (turingtest):

once again: s(t)=300000e^(-0.20)=A CONSTANT (no t's in the function) the derivative of a constant is 0 so s'(t)=0 for ALL t's including t=10 so s'(10)=0

OpenStudy (anonymous):

thats not one of the choices.. but thanks anyways

OpenStudy (turingtest):

then you wrote the problem wrong. Or it's written wrong in you textbook. you (or they) probably forgot a t somewhere. maybe it is supposed to be\[s(t)=300000e^(0.2t)\] then you get \[s'(t)=60000e^(0.2t)\]\[s'(10)=60000e^2\] is that a choice?

OpenStudy (anonymous):

no its either -2.72 -49123.85 13.62 -8120.12

OpenStudy (turingtest):

sorry I forgot the - in the exponent\[s'(t)=-60000e^(-.2t)\]\[s'(10)=-60000e^(-2)=-8120.116994\] so yes, someone forgot a "t" the answer is definitely -8120.12

OpenStudy (anonymous):

where are you getting -60000?

OpenStudy (turingtest):

chain rule on derivative of f(x)=e^(-0.2t) f'(x)=(-0.2)e^(-0.2t) so it's 300000(-0.2)

OpenStudy (anonymous):

wow thank you.

OpenStudy (turingtest):

just curious, who forgot the "t", you or your book/teacher?

OpenStudy (anonymous):

teacher. its no where on the question at all

OpenStudy (turingtest):

what a clown. let him know how badly he messed up.

OpenStudy (anonymous):

oh i will!